# Tazza: Shuffling Neural Network Parameters for Secure and Private Federated Learning

### v1

@@ 5.1. Overview of Tazza
Figure 5 illustrates the overall workflow of Tazza: 1 The federated learning server distributes a global model to each client. 2 Clients selected to participate in training share a shuffling rule via a safeguarded peer-to-peer communication. As a communication-efficient approach, Tazza simply shares the random seed for generating the shuffling order (Sec 5.2). 3 Participants of the federated learning round train the neural network model with their locally collected data. 4 Upon completing local training, each participant shuffles its local model weights with respect to the shared rule to ensure the confidentiality of the training operations (Weight Shuffling-Sec. 5.3). 5 Participating clients transmit the shuffled model parameters and a small number of input samples (as small as one) to the server. We note that the input samples are also shuffled to prevent the confidentiality of the system from being compromised. 6 With the updated/shuffled models and data samples, the server performs an integrity check via shuffled model validation on the data samples using each client’s model by measuring the similarity of resulting output vectors (Sec. 5.4). Since models and data samples are shuffled to satisfy the permutation invariance of the neural network, the resulting output vectors are still legitimate. 7 Exploiting the similarity between the clients calculated in the previous step, Tazza employs a clustering algorithm to cluster the clients. As malicious and benign models exhibit different computation results, grouping the clients based on their similarity effectively assures that benign clients are not affected by malicious model updates.
To summarize, in Step 4, the weight shuffling process ensures that by shuffling the weights, the server is not able to infer information about the client’s local data through model analysis. Nevertheless, since weights are shuffled while preserving permutation invariance, the vectors calculated in Step 6 still remain valid. This enables the shuffled validation process to effectively identify malicious clients attempting poisoning attacks by clustering them in Step 7.
@@ 5.2. Shuffling Rule Sharing
Before delving into the detailed design of Tazza, we first address a potential issue regarding clients independently shuffling their model weights in differing orders, which would void the validation process and lead to degraded aggregated model performance. Therefore, it is important for clients to agree upon a unified weight-shuffling order. However, the challenge here is that if the server gains access to this unified shuffling rule, it could potentially revert the shuffled models to their original state, posing a security risk. To mitigate this, the exchange of shuffling rules must be conducted exclusively among the clients without server involvement.
Secure peer-to-peer communication protocols, such as zero-knowledge proof (Boo et al., 2021) and Byzantine fault-tolerant methods (Castro et al., 1999), provide practical and well-established mechanisms for achieving this client-only data exchange. As noted earlier, we assume that clients leverage these secure communication methods to share shuffling rules amongst themselves, ensuring that the server remains uninformed of the shuffling specifics. This assumption supports the integrity and confidentiality of the proposed system.
Specifically, Tazza employs a Practical Byzantine Fault Tolerant (PBFT) mechanism (Castro et al., 1999) for exchanging its shuffling rules. The server selects one participating client as the head, which distributes a shuffling rule to others. Participants that receive the rule then broadcast the rule to non-participating clients to reach consensus. Even if a malicious client is chosen as the leader, PBFT ensures that consensus is achieved and the shuffling rule remains secure and consistent across all clients, despite potential attacks.
@@ 5.3. Weight Shuffling
In essence, weight shuffling in Tazza consists of two core operations: row shuffling and column shuffling. Consider a matrix multiplication between the input matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ and the weight matrix $\mathbf{W}\in\mathbb{R}^{d\times h}$, where $N$, $d$, and $h$ denote the number of vectors, input dimension, and embedding dimension, respectively. To shuffle the weight rows, a row-permutation matrix $R$ can be applied to $\mathbf{W}$ (i.e., $\mathbf{R}\times\mathbf{W}$). This operation is computationally equivalent to directly shuffling the input, thereby ensuring data confidentiality. Furthermore, by applying the same permutation to the columns of $\mathbf{X}$ via $\mathbf{R}^{\top}$ (i.e., $\mathbf{X}\times\mathbf{R}^{\top}$), the computation results remain unchanged due to the property of permutation invariance.
The weights can be shuffled further with respect to the columns by multiplying the weight matrix $\mathbf{W}$ with a column-permutation matrix $\mathbf{C}$ (i.e., $\mathbf{W}\times\mathbf{C}$). Since column shuffling affects only the column order of the output and not the computation itself, no additional permutation needs to be applied to the input. However, when considering the bias term $\mathbf{b}\in\mathbb{R}^{h}$ in a linear layer (i.e., $\mathbf{X}\times\mathbf{W}+\mathbf{b}$), the same column permutation must be applied to $\mathbf{b}$ to ensure proper alignment with the shuffled weights. Since the output of a layer serves as the input for the subsequent layer, the column shuffling order of the current layer can be directly inherited as the row shuffling order for the weights of the next layer.
This shuffling operation for matrix multiplication (i.e., linear layers) leverages permutation invariance and equivariance, which form the foundation of most neural network computations. However, simply uniformly applying this approach to all weights may not always work as intended. This is because, while the underlying principle remains consistent, specific behaviors and designs of models can vary. Using we elaborate on how this shuffling operation can be adapted for various neural network architectures in Section 6.
@@ 5.4.1. Shuffled Model Validation
Malicious models produce outputs that deviate from those of benign models, revealing distinctive patterns. Tazza identifies such deviations by computing pairwise cosine similarity among the output vectors of all models (Park et al., 2023). Malicious models typically exhibit lower cosine similarity with benign ones, providing a basis for their classification and isolation.
Clients are then clustered based on their pairwise cosine similarities, with the similarity matrix ($S$) transformed into a distance matrix ($1-S$) to interpret similarity as a distance metric, which is used as input to the DBSCAN algorithm (Ester et al., 1996). Tazza leverages DBSCAN for its ability to identify outliers and its independence from predefined cluster numbers, making it particularly well-suited for federated learning scenarios where the presence of attackers is unknown.
@@ 5.4.2. Cluster-aware Aggregation
Finally, we implement cluster-aware aggregation to fully segregate malicious models from benign ones. In each federated learning round, computed cluster labels serve as pseudo-labels, reflecting the similarities among clients within that round only. These labels are independent across rounds (e.g., labels from round $t$ have no relation to those from round $t+1$). To ensure consistency, Tazza assigns global cluster labels by leveraging clustering histories, enabling effective cross-round clustering. Clients are categorized into two groups: (1) those participating for the first time, who lack a global cluster label, and (2) those with prior participation and an assigned global cluster.
Upon performing shuffled model validation, if all clients in a cluster are first-time participants without any global cluster assignment, Tazza creates and assigns new global clusters. Conversely, if a pseudo-cluster contains clients with pre-existing global cluster labels, all clients in the cluster are assigned to the global cluster label of those clients.
Furthermore, if clients who were previously part of different global clusters are grouped in the same pseudo-cluster during the current round, Tazza merges their global clusters. However, instead of merging all clients from both clusters indiscriminately, Tazza adopts a conservative approach by reassigning only the overlapping clients to the new cluster. This precautionary strategy ensures that a single attacker infiltrating a benign cluster cannot lead to all attackers being included in the same benign cluster.
When global clusters are assigned for participating clients, Tazza performs aggregation within each cluster based on the clustering results, producing an updated model for each cluster. We note that Tazza employs averaging as the aggregation algorithm while alternative methods such as median or trimmed mean can also be applied, offering flexibility in system design. By aggregating models only within the same cluster, Tazza prevents compromised models from contaminating the global model shared across other clusters.
@@ 6. Implementation
(1) $\begin{split}h^{(0)}&=x,\\ h^{(l)}&=\sigma(W^{(l)}h^{(l-1)}+b^{(l)}),\quad l=1,\dots,L-1,\\ y&=W^{(L)}h^{(L-1)}+b^{(L)},\end{split}$
@@
A critical observation is that commonly used activation functions are parameter-free and permutation-equivariant, meaning their intrinsic behavior remains unaffected by input shuffling. As a result, activation functions can be excluded from considerations in weight shuffling operations. This leaves the linear transformations in each layer, represented as matrix multiplications ($W^{(l)}h^{(l-1)}$), as the primary focus. The weight matrices and bias vectors in these transformations can be shuffled according to the previously outlined strategies, ensuring that computational outcomes remain consistent despite weight shuffling. By applying this shuffling mechanism iteratively across layers, the entire MLP architecture retains compatibility with weight shuffling operations, preserving computational integrity while safeguarding model confidentiality.
@@ 7.1. Experiment Setup
$\bullet$ MIT-BIH Arrhythmia Database (Moody and Mark, 2001, 1992) is a widely used electrocardiogram (ECG) benchmark dataset comprising approximately 70K five-second ECG signals. We implement an RNN  (Rumelhart et al., 1986) to classify three arrhythmia types along with normal cardiac patterns. This dataset is utilized to explore the potential extension of Tazza to time-series inputs.
@@
For integrity attacks, we evaluate two baseline configurations: FedAvg (McMahan et al., 2017) and Oracle, representing naive and ideal model aggregation scenarios, respectively. Specifically, FedAvg does not employ security mechanisms, leaving it vulnerable to malicious updates, while Oracle assumes perfect knowledge of attackers, with the capability of fully excluding malicious models’ contributions on the global model.
We also evaluate robust aggregation methods leveraging statistical priors, including Median and Trimmed Mean (Yin et al., 2018). Median replaces averaging with the median to mitigate extreme outlier influence, while Trimmed Mean excludes outliers before averaging, reducing the impact of malicious updates. Additionally, we compare Tazza with MultiKrum (Blanchard et al., 2017) and FLTrust (Cao et al., 2020). MultiKrum selects updates with the smallest pairwise Euclidean distances after discarding the most distant ones. FLTrust assigns trust scores based on cosine similarity with a server-trained reference model, adjusting aggregation weights to suppress malicious updates.
@@
Miscellaneous Configurations. Throughout the evaluation, unless otherwise specified, datasets are distributed across 100 clients, consisting of 25 attackers, with data shared via a Dirichlet distribution with $\alpha$=0.5 (Hsu et al., 2019). In each federated learning round, 25 participants are selected, and their models are trained locally for 5 epochs using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 1e-3 and a batch size of 64. Besides the evaluations performed on real embedded computing platforms in Section 7.5, all experiments were conducted on a server with an NVIDIA RTX 2080 GPU, an Intel Xeon Silver 4210 CPU 2.20GHz, and 64GB RAM.
@@ 7.2. Overall Performance
Figure 8. LPIPS and accuracy of varying federated learning configurations. $\blacktriangle$ and $\bullet$ denote the applying of gradient pruning and DPSGD, respectively.
@@
Figure 8 presents the global model accuracy under label-flipping attacks and the LPIPS across various integrity and confidentiality defense schemes. The legend indicates different integrity defense schemes, with triangle ($\blacktriangle$) and circle ($\bullet$) markers representing confidentiality defense schemes using gradient pruning and DPSGD, respectively. Gradient pruning is applied with pruning ratios of 70%, 90%, and 99%, while DPSGD uses three perturbation levels: small, medium, and large, as outlined in prior work (Abadi et al., 2016). Each scheme includes three plots: the leftmost plot corresponds to the weakest confidentiality defense configuration (70% pruning for gradient pruning and small perturbation for DPSGD), while the rightmost plot represents the strongest configuration (99% pruning for gradient pruning and large perturbation for DPSGD).
Existing confidentiality defense mechanisms improve data privacy, achieving higher LPIPS values, but often result in reduced accuracy. In contrast, Tazza achieves both high LPIPS and high accuracy, demonstrating its robustness against integrity and confidentiality attacks. Notably, even under an idealized (Oracle) scenario where the server perfectly excludes malicious updates during aggregation, the model accuracy remains lower than that of Tazza, due to inaccuracies introduced by confidentiality defense mechanisms in the aggregated weights.

### v2

@@ 5.1. Overview of Tazza
Figure 7 illustrates the overall workflow of Tazza: 1 The federated learning server distributes a global model to each client. 2 Clients selected to participate in training share a shuffling rule via a safeguarded peer-to-peer communication. As a communication-efficient approach, Tazza simply shares the random seed for generating the shuffling order (Sec 5.2). 3 Participants of the federated learning round train the neural network model with their locally collected data. 4 Upon completing local training, each participant shuffles its local model weights with respect to the shared rule to ensure the confidentiality of the training operations (Weight Shuffling-Sec. 5.3). 5 Participating clients transmit the shuffled model parameters and a small number of input samples (as small as one) to the server. We note that the input samples are also shuffled to prevent the confidentiality of the system from being compromised. 6 With the updated/shuffled models and data samples, the server performs an integrity check via shuffled model validation on the data samples using each client’s model by measuring the similarity of resulting output vectors (Sec. 5.4). Since models and data samples are shuffled to satisfy the permutation invariance of the neural network based on the shared shuffling rule in 2, the resulting output vectors are still legitimate. 7 Exploiting the similarity between the clients calculated in the previous step, Tazza employs a clustering algorithm to cluster the clients. As malicious and benign models exhibit different computation results, grouping the clients based on their similarity effectively assures that benign clients are not affected by malicious model updates.
To summarize, in Step 4, the weight shuffling process ensures that by shuffling the weights, the server is not able to infer information about the client’s local data through model analysis. Nevertheless, since weights are shuffled while preserving permutation invariance, the vectors calculated in Step 6 still remain valid. This enables the shuffled validation process to effectively identify malicious clients attempting poisoning attacks by clustering them in Step 7. In addition, despite the altered order, model aggregation can take place seamlessly as all clients are employing the same shuffling rule.
@@ 5.2. Shuffling Rule Sharing
Before delving into the detailed design of Tazza, we first address a potential issue regarding clients independently shuffling their model weights in differing orders, which would void the validation process and lead to degraded aggregated model performance. Therefore, it is important for clients to agree upon a unified weight-shuffling order. However, the challenge here is that if the server gains access to this unified shuffling rule, it could potentially revert the shuffled models to their original state, posing a security risk. To mitigate this, the core requirement for securing the system lies in establishing an information asymmetry between the server and clients regarding the shuffling rule. To achieve this, existing security mechanisms such as secure multi-party computation (Zhao et al., 2019), zero-knowledge proofs (Boo et al., 2021), confidential computing (Mo et al., 2021), or peer-to-peer communication (Castro et al., 1999) can be employed to fulfill this prerequisite. As previously mentioned, we assume that clients utilize these secure communication methods to exchange shuffling rules among themselves, ensuring that the server remains unaware of the specific shuffling details.
Specifically, Tazza employs a Practical Byzantine Fault Tolerant (PBFT) mechanism (Castro et al., 1999) for exchanging its shuffling rules. The server selects one participating client as the head, which distributes a shuffling rule to others. Participants who receive the rule then broadcast the rule to non-participating clients to reach a consensus. Even if a malicious client is chosen as the leader, PBFT ensures that consensus is achieved and the shuffling rule remains secure and consistent across all clients, despite potential attacks.
In detail, during each federated training round, clients exchange two shuffling rules to support model updates. First, the model distributed by the server is encrypted using the shuffling rule from the previous round, which clients retrieve through secure communication from clients that have participated in the specific round. Additionally, clients share a new shuffling rule among the current round’s participants to encrypt the locally updated model. As previously noted, the shared rules are lightweight, being as simple as a single random seed to minimize the communication overhead.
@@ 5.3. Weight Shuffling
In essence, weight shuffling in Tazza consists of two core operations: row shuffling and column shuffling. Consider a matrix multiplication between the input matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ and the weight matrix $\mathbf{W}\in\mathbb{R}^{d\times h}$, where $N$, $d$, and $h$ denote the number of vectors, input dimension, and embedding dimension, respectively. To shuffle the weight rows, a row-permutation matrix $\mathbf{R}$ can be applied to $\mathbf{W}$ (i.e., $\mathbf{R}\times\mathbf{W}$). This operation is computationally equivalent to directly shuffling the input, thereby ensuring data confidentiality. Here, we note that a permutation matrix $\mathbf{R}$ satisfies the following lemma: $\mathbf{R}^{\top}\mathbf{R}=\mathbf{I}$ where $\mathbf{I}$ denotes the identity matrix. Based on this, by applying the same permutation to $\mathbf{X}$ via $\mathbf{R}^{\top}$ (i.e., $\mathbf{X}\times\mathbf{R}^{\top}$), the computation results remain unchanged due to the property of permutation invariance, since $(\mathbf{X}\mathbf{R}^{\top})(\mathbf{R}\mathbf{W})=\mathbf{XIW}=\mathbf{XW}$ (c.f., Figure 3 (b)).
The weights can be shuffled further with respect to the columns by multiplying the weight matrix $\mathbf{W}$ with a column-permutation matrix $\mathbf{C}$ (i.e., $\mathbf{W}\times\mathbf{C}$). Since column shuffling affects only the column order of the output and not the computation itself, no additional permutation needs to be applied to the input (c.f., Figure 3 (c)). However, when considering the bias term $\mathbf{b}\in\mathbb{R}^{h}$ in a linear layer (i.e., $\mathbf{X}\times\mathbf{W}+\mathbf{b}$), the same column permutation must be applied to $\mathbf{b}$ to ensure proper alignment with the shuffled weights. Since the output of a layer serves as the input for the subsequent layer, the column shuffling order of the current layer can be directly inherited as the row shuffling order for the weights of the next layer.
This shuffling operation for matrix multiplication (i.e., linear layers) leverages permutation invariance and equivariance, which form the foundation of most neural network computations. However, simply uniformly applying this approach to all weights may not always work as intended. This is because, while the underlying principle remains consistent, specific behaviors and designs of models can vary. We elaborate on how this shuffling operation can be adapted for various complex and advanced neural network architectures in Section 6.
@@ 5.4.1. Shuffled Model Validation
Malicious models produce outputs that deviate from those of benign models, revealing distinctive patterns. Tazza identifies such deviations by computing pairwise cosine similarity among the output vectors of all models (Park et al., 2023a). Malicious models typically exhibit lower cosine similarity with benign ones, providing a basis for their classification and isolation. Clients are then clustered based on their pairwise cosine similarities, with the similarity matrix ($S$) transformed into a distance matrix ($1-S$) to interpret similarity as a distance metric, which is used as input to the DBSCAN algorithm (Ester et al., 1996). Tazza leverages DBSCAN for its ability to identify outliers and its independence from predefined cluster numbers, making it particularly well-suited for federated learning scenarios where the presence of attackers is unknown.
@@ 5.4.2. Cluster-aware Aggregation
Input: Client models and pseudo-cluster labels for round $t$, global cluster labels $G$
Output: Updated global cluster labels and aggregated models per cluster
1
2Initialize: New global cluster set $G_{\text{new}}\leftarrow G$;
3
4foreach pseudo-cluster $C_{p}$ in round $t$ do
5       Classify clients in $C_{p}$:
6        $C_{\text{new}}\leftarrow$ clients without global labels;
7        $C_{\text{existing}}\leftarrow$ clients with global labels;
8
9      if $C_{\text{existing}}=\emptyset$ then
10             Assign a new global cluster $g_{\text{new}}$ to all clients in $C_{p}$;
11             Update $G_{\text{new}}\leftarrow G_{\text{new}}\cup\{g_{\text{new}}\}$;
12
13      else
14             Identify the dominant global label $g_{\text{dom}}$ in $C_{\text{existing}}$;
15             Assign $g_{\text{dom}}$ to all clients in $C_{p}$;
16
17            if clients with different global labels are grouped in $C_{p}$ then
18                   Reassign overlapping clients in $C_{p}$ to a new merged cluster;
19                   Update $G_{\text{new}}$ with the merged cluster;
20
21
22
23foreach global cluster $g$ in $G_{\text{new}}$ do
24       Perform model aggregation for clients in $g$;
25
26return Updated global clusters $G_{\text{new}}$ and aggregated models;
Algorithm 1 Cluster-Aware Aggregation in Tazza
Finally, we implement cluster-aware aggregation to fully segregate malicious models from benign ones. Algorithm 1 presents the pseudo-code for the cluster-aware aggregation process. In each federated learning round, computed cluster labels from DBSCAN serve as pseudo-labels, reflecting the similarities among clients within that round only. These labels are independent across rounds (e.g., labels from round $t$ have no relation to those from round $t+1$). To ensure consistency, Tazza assigns global cluster labels by leveraging clustering histories, enabling effective cross-round clustering. Clients are categorized into two groups: (1) those participating for the first time, who lack a global cluster label, and (2) those with prior participation and an assigned global cluster (line 3-5 in Algorithm 1).
Upon performing shuffled model validation, if all clients in a cluster are first-time participants without any global cluster assignment, Tazza creates and assigns new global clusters (line 6-8 in Algorithm 1). Conversely, if a pseudo-cluster contains clients with pre-existing global cluster labels, all clients in the cluster are assigned to the global cluster label of those clients (line 10-11 in Algorithm 1).
Furthermore, if clients who were previously part of different global clusters are grouped in the same pseudo-cluster during the current round, Tazza merges their global clusters. However, instead of merging all clients from both clusters indiscriminately, Tazza adopts a conservative approach by reassigning only the overlapping clients to the new cluster (line 12-14 in Algorithm 1). This precautionary strategy ensures that a single attacker infiltrating a benign cluster cannot lead to all attackers being included in the same benign cluster.
When global clusters are assigned for participating clients, Tazza performs aggregation within each cluster based on the clustering results, producing an updated model for each cluster (line 15-16 in Algorithm 1). We note that Tazza employs averaging as the aggregation algorithm while alternative methods such as median or trimmed mean can also be applied, offering flexibility in system design. By aggregating models only within the same cluster, Tazza prevents compromised models from contaminating the global model shared across other clusters. We also point out that, although the server is unaware of how model parameters are shuffled, it can still seamlessly aggregate these parameters since they are organized according to consistent rules as discussed in Section 5.3.
@@ 6. Implementation
def shuffle_linear(fc:nn.Module, shuffle_rule:int):
    Pseudo-code for linear layer (MLP) weight shuffling
    fc = (input_dim, hidden_dim) shaped linear layer
    (i.e., nn.Linear(input_dim, hidden_dim))
    # fix random seed to generate unified shuffling indices across clients
    torch.manual_seed(shuffle_rule)
    hidden_dim, in_dim = fc.weight.size()
    in_shuffle_idx = torch.randperm(in_dim)
    # unshuffle_idx = torch.argsort(in_shuffle_idx) # to unshuffle
    hidden_shuffle_idx = torch.randperm(hidden_dim)
    fc.weight = torch.nn.Parameter(fc.weight[hidden_shuffle_idx][:, in_shuffle_idx])
    fc.bias   = torch.nn.Parameter(fc.bias[hidden_shuffle_idx])
    return fc
List of Listings 1 PyTorch implementation of the weight shuffling module for a linear layer.
@@
(1) $\begin{split}h^{(0)}&=x,\\ h^{(l)}&=\sigma(h^{(l-1)}W^{(l)\top}+b^{(l)}),\quad l=1,\dots,L-1,\\ y&=h^{(L-1)}W^{(L)\top}+b^{(L)},\end{split}$
@@
A critical observation is that commonly used activation functions are parameter-free and permutation-equivariant, meaning their intrinsic behavior remains unaffected by input shuffling. As a result, activation functions can be excluded from considerations in weight shuffling operations. This leaves the linear transformations in each layer, represented as matrix multiplications ($h^{(l-1)}W^{(l)\top}$), as the primary focus. The weight matrices and bias vectors in these transformations can be shuffled according to the previously outlined strategies, ensuring that computational outcomes remain consistent despite weight shuffling. By applying this shuffling mechanism iteratively across layers, the entire MLP architecture retains compatibility with weight shuffling operations, preserving computational integrity while safeguarding model confidentiality.
Listing 1 presents the PyTorch (Paszke et al., 2019) implementation of weight shuffling for the affine linear transformation layer, defined as $f(\mathbf{x})=\mathbf{xW^{\top}}+\mathbf{b}$. To minimize both memory and computational overhead, particularly given the resource-constrained nature of mobile and embedded devices deployed in ubiquitous applications, we replace explicit permutation matrix multiplication with index permutation. Notably, since the linear transformation layer relies on fundamental operations, such as matrix multiplication and addition, that are widely used in neural network architectures, this approach can be easily extended to more complex and advanced architectures without loss of generality.
@@ 7.1. Experiment Setup
$\bullet$ MIT-BIH Arrhythmia Database (Moody and Mark, 2001, 1992) is a widely used electrocardiogram (ECG) benchmark dataset comprising approximately 70K five-second ECG signals. We implement an RNN  (Rumelhart et al., 1986) to classify three arrhythmia types along with normal cardiac patterns. This dataset is utilized to explore the potential extension of Tazza to mobile/embedded applications that deal with time-series inputs (Park et al., 2023b; Lee et al., 2024a).
@@
For integrity attacks, we use FedAvg (McMahan et al., 2017) as a baseline, representing a naive model aggregation scenario. Specifically, FedAvg does not employ security mechanisms, leaving it vulnerable to malicious updates. We also evaluate robust aggregation methods leveraging statistical priors, including Median and Trimmed Mean (Yin et al., 2018). Median replaces averaging with the median to mitigate extreme outlier influence, while Trimmed Mean excludes outliers before averaging, reducing the impact of malicious updates. Additionally, we compare Tazza with MultiKrum (Blanchard et al., 2017) and FLTrust (Cao et al., 2020), which are representative frameworks for similarity based approaches. MultiKrum selects updates with the smallest pairwise Euclidean distances after discarding the most distant ones. FLTrust assigns trust scores based on cosine similarity with a server-trained reference model, adjusting aggregation weights to suppress malicious updates.
@@
Miscellaneous Configurations. Throughout the evaluation, unless otherwise specified, datasets are distributed across 100 clients, consisting of 25 attackers, with data shared via a Dirichlet distribution with $\alpha$=0.5 (Hsu et al., 2019). In each federated learning round, 10 participants are selected, and their models are trained locally for 5 epochs using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 1e-3 and a batch size of 64. Besides the evaluations performed on real embedded computing platforms in Section 7.5, all experiments were conducted on a server with an NVIDIA RTX 2080 GPU, an Intel Xeon Silver 4210 CPU 2.20GHz, and 64GB RAM.
@@ 7.2. Overall Performance
Figure 10. LPIPS and accuracy of varying federated learning configurations. $\blacktriangle$ and $\bullet$ denote applying gradient pruning and DPSGD, respectively.
@@
Figure 10 presents the global model accuracy under label-flipping attacks and the LPIPS across various integrity and confidentiality defense schemes. Additionally, since each cluster in Tazza has distinct model parameters, we report the average accuracy of the global model on the global test dataset across clusters that contain at least one benign client. The legend indicates different integrity defense schemes, with triangle ($\blacktriangle$) and circle ($\bullet$) markers representing confidentiality defense schemes using gradient pruning and DPSGD, respectively. Gradient pruning is applied with pruning ratios of 70%, 90%, and 99%, while DPSGD uses three perturbation levels: small, medium, and large, as outlined in prior work (Abadi et al., 2016). Each scheme includes three plots: the leftmost plot corresponds to the weakest confidentiality defense configuration (70% pruning for gradient pruning and small perturbation for DPSGD), while the rightmost plot represents the strongest configuration (99% pruning for gradient pruning and large perturbation for DPSGD).
Existing confidentiality defense mechanisms improve data privacy, achieving higher LPIPS values, but often result in reduced accuracy. In contrast, Tazza achieves both high LPIPS and high accuracy, demonstrating its robustness against integrity and confidentiality attacks.